OpenStudy (anonymous):

a ball is thrown straight up into the air with an initial speed of 18.2 m/s. How far has it gone after 1.00s?

OpenStudy (mrnood):

s= ut + 0.5 at^2 s= distance u= initial speed (= +18.2m/s) a= -9.81m/s^2 t= 1s

OpenStudy (mrnood):

Be careful with the signs Initial speed is in opposite direction to acceleration

OpenStudy (anonymous):

i thought a= -9.80? @MrNood

OpenStudy (mrnood):

it really doesn't matter in this case ( the difference you are talking about is 0.1%) I believe -9.81 is a more accurate figure. However - use whatever value you want - but put it into the equation above and solve for s

OpenStudy (anonymous):

i got 66.2 but it said wrong @MrNood

OpenStudy (mrnood):

to quote from Wikipedia: The precise strength of Earth's gravity varies depending on location. The nominal "average" value at the Earth's surface, known as standard gravity is, by definition,[2] 9.80665 m/s2

OpenStudy (mrnood):

That's cos it IS wrong. I have given you the equation and all the values. If you show your working I will help you see where you made the mistake, but I will not post the answer

OpenStudy (mrnood):

start by calculating \[\frac{at ^{2}}{ 2 } \] using the values I gave above

OpenStudy (mrnood):

gone offline as soon as I said I wouldn't give answer